OpenStudy (anonymous):

An object with a mass of 8.6 kg experiences a force of 12.9 N. What is the acceleration of the object?

OpenStudy (anonymous):

f=ma 12.9=8.6*a a=12.9/8.6=1.5 m/(s)square